# Transparent Networks for Multivariate Time Series

### v2

@@ 5.1 Experimental Setup
Baselines: We compare our GATSM with 12 baselines, which can be categorized into four groups: 1) Black-box tabular models include extreme gradient boosting (XGBoost) [35] and MLP. 2) Black-box time series models include simple recurrent neural network (RNN), gated recurrent unit (GRU), long short-term memory (LSTM), and Transformer [16]. 3) Transparent tabular models are simple linear model (Linear), explainable boosting machine (EBM) [23], NAM [7], NodeGAM [8], and NBM [9]. 4) NATM [30] is a transparent time series model.
Implementation: We implement XGBoost and EBM models using the xgboost and interpretml libraries, respectively. For NodeGAM, we employ the official implementation provided by its authors [8]. The remaining models are developed using PyTorch [36]. All models undergo hyperparameter tuning via Optuna [37]. The pytorch-based models are optimized with the Adam with decoupled weight decay (AdamW) [38] optimizer on an NVIDIA A100 GPU. Model training is halted if the validation loss does not decrease over 20 epochs. We use mean squared error for the forecasting tasks, and for classification tasks, we use cross-entropy loss. Further details of the model implementations and hyper-parameters are provided in Appendix C.
5.2 Comparison with baselines
Table 3: Predictive performance comparison of various models.
[…]
Black-box Tabular Model XGBoost 0.094 0.002 0.532 0.679 0.707 0.816 0.424 0.200 8.500
($\pm$0.137) ($\pm$0.002) ($\pm$0.019) ($\pm$0.094) ($\pm$0.015) ($\pm$0.007) ($\pm$0.012) ($\pm$0.049) ($\pm$4.000)
MLP 0.459 0.011 0.423 0.654 0.842 0.786 0.417 0.211 7.375
($\pm$0.101) ($\pm$0.004) ($\pm$0.031) ($\pm$0.082) ($\pm$0.014) ($\pm$0.007) ($\pm$0.008) ($\pm$0.065) ($\pm$2.134)
Black-box Time Series Model RNN 0.320 0.068 0.644 0.661 0.581 0.782 0.422 0.592 7.750
($\pm$0.122) ($\pm$0.020) ($\pm$0.032) ($\pm$0.078) ($\pm$0.040) ($\pm$0.009) ($\pm$0.029) ($\pm$0.110) ($\pm$2.712)
GRU 0.435 0.089 0.701 0.694 0.818 0.785 0.629 0.931 4.375
($\pm$0.107) ($\pm$0.034) ($\pm$0.018) ($\pm$0.052) ($\pm$0.014) ($\pm$0.010) ($\pm$0.013) ($\pm$0.045) ($\pm$2.669)
LSTM 0.359 0.090 0.683 0.648 0.790 0.779 0.491 0.908 6.375
($\pm$0.112) ($\pm$0.031) ($\pm$0.026) ($\pm$0.042) ($\pm$0.020) ($\pm$0.008) ($\pm$0.082) ($\pm$0.035) ($\pm$3.623)
Transformer 0.263 0.098 0.711 0.690 0.844 0.789 0.679 0.967 4.000
($\pm$0.263) ($\pm$0.035) ($\pm$0.027) ($\pm$0.040) ($\pm$0.019) ($\pm$0.010) ($\pm$0.019) ($\pm$0.029) ($\pm$3.703)
Transparent Tabular Model Linear 0.482 0.004 0.241 0.637 0.838 0.723 0.311 0.206 10.125
($\pm$0.112) ($\pm$0.001) ($\pm$0.019) ($\pm$0.070) ($\pm$0.017) ($\pm$0.011) ($\pm$0.010) ($\pm$0.045) ($\pm$3.871)
EBM -0.200 0.004 0.324 0.666 0.729 0.802 0.408 0.164 9.750
($\pm$0.409) ($\pm$0.001) ($\pm$0.014) ($\pm$0.056) ($\pm$0.017) ($\pm$0.011) ($\pm$0.016) ($\pm$0.053) ($\pm$3.284)
NAM 0.363 0.006 0.300 0.645 0.853 0.800 0.400 0.242 7.875
($\pm$0.218) ($\pm$0.002) ($\pm$0.013) ($\pm$0.026) ($\pm$0.014) ($\pm$0.006) ($\pm$0.011) ($\pm$0.040) ($\pm$3.643)
NodeGAM 0.398 0.006 0.380 0.681 0.854 0.802 0.400 0.247 6.375
($\pm$0.195) ($\pm$0.002) ($\pm$0.032) ($\pm$0.046) ($\pm$0.013) ($\pm$0.007) ($\pm$0.028) ($\pm$0.012) ($\pm$3.623)
NBM 0.330 0.007 0.301 0.716 0.852 0.799 0.388 0.189 7.875
($\pm$0.251) ($\pm$0.003) ($\pm$0.012) ($\pm$0.039) ($\pm$0.014) ($\pm$0.006) ($\pm$0.014) ($\pm$0.029) ($\pm$3.603)
Transparent Time Series Model NATM 0.304 0.038 0.548 0.724 N/A N/A 0.452 0.878 5.667
($\pm$0.122) ($\pm$0.011) ($\pm$0.028) ($\pm$0.043) ($\pm$0.010) ($\pm$0.058) ($\pm$2.582)
GATSM (ours) 0.493 0.073 0.583 0.843 0.853 0.797 0.570 0.956 3.125
($\pm$0.173) ($\pm$0.027) ($\pm$0.026) ($\pm$0.025) ($\pm$0.015) ($\pm$0.007) ($\pm$0.024) ($\pm$0.027) ($\pm$1.808)
Table 3 shows the predictive performances of the experimental models. We report mean scores and standard deviations over five different random seeds. For the forecasting datasets, we evaluate $R^{2}$ scores. For the binary classification datasets, we assess the area under the receiver operating characteristic curve (AUROC). For the multi-class classification datasets, we measure accuracy. We highlight the best-performing model in bold and underline the second-best model. Since the tabular models cannot handle time series, they only take $\textbf{x}_{t}$ to produce $y_{t}$.
On the Energy and Heartbeat datasets, which are small in size, our GATSM demonstrates the best performance, indicating strong generalization ability. EBM, XGBoost, and Transformer struggle with overfitting on the Energy dataset. For the Mortality and Sepsis datasets, there is no significant performance difference between tabular and time series models, nor between black-box and transparent models. This suggests that these two healthcare datasets lack significant temporal patterns and feature interactions. It is likely that seasonal patterns are hard to detect in medical data, and the patient’s current condition already encapsulates previous conditions, making historical data less crucial. Since these datasets contain variable-length time series, the performance of NATM, which can only handle fixed-length time series, is not available. On the Rainfall, AirQuality, LSST, and NATOPS datasets, the time series models significantly outperform the tabular models, indicating that these datasets contain important temporal patterns that tabular models cannot capture. Additionally, the black-box models outperform the transparent models, suggesting that these datasets have higher-order feature interactions that transparent models cannot capture. Nevertheless, GATSM is the best model within the transparent model group and performs comparably to Transformer. Overall, GATSM achieved the best average rank in the experiments, followed by the Transformer, indicating GATSM’s superiority. Additional experiments on model throughput and an ablation study on the basis functions are presented in Appendix D.
5.3 Ablation study
Choice of feature function: We evaluate the performance of GATSM by changing the feature functions using three models: Linear, NAM, and NBM. Table 4 presents the results of this experiment. The simple linear function performs poorly because it lacks the capability to capture non-linear relationships. In contrast, NAM, which can capture non-linearity, shows improved performance over the linear function. However, NBM stands out by achieving the best performance in six out of eight datasets. This indicates that the basis strategy of NBM is highly effective for time series data.
Table 4: Ablation study on different feature functions.
Design of temporal module: We evaluate the performance of GATSM by modifying the design of the temporal module. The results are presented in Table 5. GATSM without the temporal module (Base) fails to learn temporal patterns and shows poor performance in the experiment. GATSM with only positional encoding (Base + PE) also shows similar performance to the Base, indicating that positional encoding alone is insufficient for capturing effective temporal patterns. GATSM with only multi-head attention (Base + MHA) outperforms the previous two methods, demonstrating that the MHA mechanism is beneficial for capturing temporal patterns. Finally, our full GATSM (Base + PE + MHA) significantly outperforms the other methods, suggesting that the combination of PE and MHA creates a synergistic effect. Consistent with our previous findings in section 5.2, all four methods show similar performances on the Mortality and Sepsis datasets, which lack significant temporal patterns.
Table 5: Ablation study on the temporal module.
5.4 Interpretation
In this section, we visualize four interpretations of GATSM’s predictions on the AirQuality dataset. In addition, interpretations for the Rainfall dataset can be found in Appendix E.
Figure 4: Local time-independent feature contributions.
Figure 5: Local time-dependent feature contributions.
Time-step importance: We plot the average attention scores at the last time step $T$ in Figure 3. The process for extracting the average attention score of time step $u$ at time step $t$ is formalized as $\sum_{k=1}^{K}a_{k,t,u}$. This process is repeated over all data samples, and the results are averaged. Based on Figure 3, it seems that GATSM pays more attention to the initial and last states than to the intermediate states. This indicates that the current concentration of particulate matter depends on the initial state.
Global feature contribution: Figure 3 illustrates the global behavior of features in the AirQuality dataset, with red bars indicating the density of training samples. We extract $\sum_{k=1}^{K}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ from GATSM and repeat this process over the range of minimum to maximum feature values to plot the line. We found that the behavior of SO2, O3, and windspeed is inconsistent with prior human knowledge. Typically, high levels of SO2 and O3 are associated with poor air quality. However, GATSM learned that particulate matter concentration starts to decrease when SO2 exceeds 10 and O3 exceeds 5. This discrepancy may be due to sparse training samples in these regions, leading to insufficient training, or there may be interactions with other features. Another known fact is that high windspeed decreases particulate matter concentration. This is consistent when windspeed is below 0.7 in our observation. However, particulate matter concentration drastically increases when windspeed exceeds 0.7, likely due to the wind causing yellow dust.
Local time-independent feature contribution: To interpret the prediction of a data sample, we plot the local time-independent feature contributions, $\sum_{k=1}^{K}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$, in Figure 4. The main x-axis (blue) represents feature contribution, the sub x-axis (red) represents feature value, and the y-axis represents time steps. We found that SO2, NO2, CO, and O3 have positive correlations. In contrast, temperature, pressure, dew point, and windspeed have negative correlations. These are consistent with the global interpretations shown in Figure 3. Rainfall has the same values across all time steps.
Local time-dependent feature contribution: We also visualize the local time-dependent feature contributions, $\sum_{k=1}^{K}a_{k,t,u}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$. Figure 5 illustrates the interpretation of the same data sample as in Figure 4. The time-dependent interpretation differs slightly from the time-independent interpretation. We found that there are time lags in SO2, NO2, CO, and O3, meaning previous feature values affect current feature contributions. For example, in the case of SO2, low feature values around time step 5 lead to low feature contributions around time step 13.
6 Future Works & Conclusion
Although GATSM achieved state-of-the-art performance within the transparent model category, it has several limitations. This section discusses these limitations and suggests future work to address them. GAMs have relatively slower computational times and larger model sizes compared to black-box models because they require the same number of feature functions as input features. To address this problem, methods such as the basis strategy can be proposed to reduce the number of feature functions, or entirely new methods for transparent models can be developed. The attention mechanism in GATSM may be a bottleneck. Fast attention mechanisms proposed in the literature [39, 40, 41, 42, 43], or the recently proposed Mamba [44], can help overcome this limitation. Existing time series models, including GATSM, only handle discrete time series and have limited length generalization ability, resulting in significantly reduced performance when very long sequences, unseen during training, are input. Extending GATSM to continuous models using NeuralODE [45] or HiPPO [46] could address this issue. GATSM still cannot learn higher-order feature interactions internally and shows low performance on complex datasets. Feature interaction methods proposed for transparent models may help address this problem [15, 29].
In this papre, we proposed a novel transparent model for time series named GATSM. GATSM consists of time-sharing NBM and the temporal module to effectively learn feature representations and temporal patterns while maintaining transparency. The experimental results demonstrated that GATSM has superior generalization ability and is the only transparent model with performance comparable to Transformer. We provided various visual interpretations of GATSM, demonstrated that GATSM capture interesting patterns in time series data. We anticipate that GATSM will be widely adopted in various fields and demonstrate strong performance. The broader impacts of GATSM across various fields can be found in Appendix A.
Appendix A Broader impact
We discuss the expected impacts of GATSM across various fields.
  • Scientific discovery: Transparent models have already been used in various research fields for scientific discovery [47, 48]. GATSM also can be applied to these domains to obtain novel scientific insights.
Appendix B Dataset details
@@
We use eight publicly available datasets for our experiments. Three datasets - Energy, Rainfall, and AirQuality - can be downloaded from the Monash repository [32]. Another three datasets - Heartbeat, LSST, and NATOPS - are available from the UCR repository [33]. The remaining two datasets can be downloaded from the PhysioNet [34]. Details of the datasets are provided below:
@@
  • Energy [49]: This dataset consists of 24 features related to temperature and humidity from sensors and weather conditions. These features are measured every 10 minutes. The goal of this dataset is to predict total energy usage.
@@
  • Rainfall [50]: This dataset consists of temperatures measured hourly. The goal of this dataset is to predict total daily rainfall in Australia.
@@
  • AirQuality [51]: This dataset consists of features related to air pollutants and meteorological data. The goal of this dataset is to predict the PM10 level in Beijing.
@@
  • Heartbeat [52]: This dataset consists of heart sounds collected from various locations on the body. Each sound was truncated to five seconds, and a spectrogram of each instance was created with a window size of 0.061 seconds with a 70% overlap. The goal of this dataset is to classify the sounds as either normal or abnormal.
@@
  • Mortality [53] This dataset consists of records of adult patients admitted to the ICU. The input features include the patient demographics, vital signs, and lab results. The goal of this dataset is to predict the in-hospital death of patients.
@@
  • Sepsis [54]: This dataset consists of records of ICU patients. The input features include patient demographics, vital signs, and lab results. The goal of this dataset is to predict sepsis six hours in advance at every time step.
@@
  • LSST [55]: This challenge dataset aims to classify astronomical time series. These time series consist of six different light curves, simulated based on the data expected from the Large Synoptic Survey Telescope (LSST).
@@
  • NATOPS [56]: This dataset aims to classify the Naval Air Training and Operating Procedures Standardization (NATOPS) motions used to control aircraft movements. It consists of 24 features representing the x, y, and z coordinates for each of the eight sensor locations attached to the body.
@@
We used get_UCR_data() and get_Monash_regression_data() functions in the tsai library [57] to load the UCR and Monash datasets.
Table 6: Optimal hyper-parameters for GATSM.
@@
Appendix C Implementation details
We use 13 models, including GATSM, for our experiments. We implement XGBoost and EBM using the xgboost [35] and interpretml [23] libraries, respectively. For NodeGAM, we employ the official implementation provided by its authors [8]. The remaining models are developed using PyTorch [36]. In addition, we implement the feature functions in NAM and NBM using grouped convolutions [58, 59] to enhance their efficiency. XGBoost and EBM are trained on two AMD EPYC 7513 CPUs, while the other models are trained on an NVIDIA A100 GPU with 80GB VRAM. All models undergo hyperparameter tuning via Optuna [37] with the Tree-structured Parzen Estimator (TPE) algorithm [60] in 100 trials. The hyperparameter search space and the optimal hyperparameters for the models are provided below:
@@ Appendix C Implementation details
  • GATSM: We tune nbm_batchnorm in the descret set {False, True}, nbm_dropout in the continuous interval [0, 0.9], attn_emb_size in the integer interval [8, 128], attn_n_heads in the integer interval [1, 8], attn_dropout in the continuous interval [0, 0.9], learning_rate in the continuous interval [1e-3, 1e-2], and weight_decay in the continuous interval [1e-6, 1e-1] on a log scale. The optimal hyper-parameters for GATSM across all experimental datasets are provided in Table 6.
Appendix D Additional experiments
D.1 Multi-step forecasting
Table 7: Statistics of the experimental datasets.
@@ D.1 Multi-step forecasting
To evaluate the performance of GATSM on the multi-step forecasting task, we conducted experiments using two sophisticated forecasting models, DLinear[61] and PatchTST[62], on three datasets: Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, Traffic [^2] http://pems.dot.ca.gov, and Weather [^3] https://www.bgc-jena.mpg.de/wetter/. Table 7 provides statistics on the experimental datasets. The models use 672 hours (four weeks) of data as input to generate 24, 48, and 72 hours-ahead predictions. Table 8 presents the mean average percentage errors (MAPEs) and standard deviations of the experimental models over five different random seeds. Overall, PatchTST demonstrates the highest predictive accuracy, while GATSM shows lower predictive performance compared to the two state-of-the-art black-box models, DLinear and PatchTST. However, our GATSM is a transparent model and offers a unique advantage in terms of interpretability.
Table 8: Predictive performance comparison on the multi-step ahead forecasting datasets.
@@
D.2 Experiment on synthetic dataset
@@
Model Cancer ($R^{2}$$\uparrow$)
DLinear 0.956
($\pm$0.009)
PatchTST 0.953
($\pm$0.006)
GATSM 0.906
($\pm$0.016)
@@
We conducted an experiment using the synthetic tumor growth dataset developed by [63], which simulates temporal changes in tumor size. This dataset is beneficial for verifying that the trained GATSM appropriately captures the influence of input features. Tumor size at any given time step is determined by three factors: prior tumor size, chemotherapy, and radiotherapy. For this experiment, we set the coefficient of chemotherapy to 10 and the coefficient of radiotherapy to 1; that is, chemotherapy more significantly reduces tumor size than radiotherapy. The experimental results in Table 9 showed that although GATSM underperforms compared to the complex black-box forecasting models, it nonetheless achieves strong predictive accuracy ($R^{2}$ > 0.9). Figure 7 and 7 illustrate the importance of time steps, and the contributions of the three factors highlight that recent time steps are notably more influential than earlier time steps due to the direct impact of prior tumor size. GATSM also effectively captured the effects of chemotherapy and radiotherapy on reducing tumor size, with chemotherapy having a significantly greater impact than radiotherapy (the scale of the y-axis in chemotherapy is greater than radiotherapy).
D.3 Inference speed
The inference speed of machine learning models is a crucial metric for real-world systems. We evaluate the throughput of various models. The results are presented in Table 10. Since the datasets have fewer features than the number of basis functions in NBM, NAM achieves higher throughput than NBM. Transparent tabular models typically exhibit fast speeds. However, their throughput significantly decreases in datasets with many features, such as Heartbeat, Mortality, and Sepsis, because they require the same number of feature functions as the number of input features. Transformer shows higher throughput than the transparent time series models because it does not require feature functions, which are the main bottleneck of transparent models. Additionally, the PyTorch implementation of Transformer uses the flash attention mechanism [64] to enhance its efficiency. NATM has slightly higher throughput than GATSM, as it does not require the attention mechanism and has fewer feature functions compared to the number of basis functions in GATSM.
Table 10: Inference throughput of different models.
Energy Rainfall AirQuality Heartbeat Mortality Sepsis LSST NATOPS
NAM 65.3K 1.8M 5.1M 139.1K 772.2K 23.9K 2.3M 147.9K
NBM 45.5K 1.1M 1.0M 55.9K 375.8K 6.5K 1.6M 85.6K
Transformer 30.9K 240.5K 174.2K 15.7K 161.9K 134.6K 214.4K 68.3K
NATM 5.3K 699.3K 241.3K 1.3K N/A N/A 28.6K 19.2K
GATSM 6.1K 350.6K 192.8K 1.2K 4.9K 3.8K 126.5K 12.5K
D.4 Number of basis functions
We evaluate GATSM by varying the number of basis functions in the time-sharing NBM. The results for forecasting, binary classification, and multi-class classification datasets are presented in Figure 8. For the Sepsis dataset, using 200 and 300 basis functions causes the out-of-memory error. For the Energy and Heartbeat datasets, performance improves up to 100 basis functions but shows no further benefit when the number of bases exceeds 100. In other datasets, performance changes are not significant with different numbers of basis functions. In addition, there is a trade-off between the number of basis functions and computational speed. Therefore, we recommend generally setting the number of basis functions to 100. Note that the performance of GATSM with this hyper-parameter depends on the dataset size and complexity. Hence, a larger number of basis functions may benefit more complex datasets.
(a) Forecasting
(b) Binary
(c) Multi-class
Appendix E Additional visualizations
In addition to the interpretations on the AirQuality dataset in section 5.4, we present another interesting interpretations of GATSM on the Rainfall dataset.

### v3

@@ Appendix B Broader impact
Table 8: Dataset statistics.
Appendix C Related Works
Various XAI studies have been conducted over the past decade (rishabh_agarwal_neural_2021; chun-hao_chang_node-gam_2022; filip_radenovic_neural_2022; sebastian_bach_pixel-wise_2015; shrikumar_learning_2017). However, they are less relevant to the transparent model that this study focuses on. Therefore, we refer readers to (ali_explainable_2023; hassija_interpreting_2024) for more detailed information on recent XAI research. In this section, we review existing transparent models closely related to our GATSM and discuss their limitations. Originally, GAMs were fitted via the backfitting algorithm using smooth splines (hastie_generalized_1986; wahba_spline_1990). Later, (yin_lou_intelligible_2012) and (nori_interpretml_2019) proposed boosted decision tree-based GAMs, which use boosted decision trees as feature functions. Spline- and tree-based GAMs have less flexibility and scalability. Thus, extending them to transfer or multi-task learning is challenging. To overcome this problem, various neural network-based GAMs have been proposed in recent years. (potts_generalized_1999) introduced generalized additive neural network, which employs 2-layer neural networks as feature functions. Similarly, (rishabh_agarwal_neural_2021) proposed neural additive model (NAM) that employs multi-layer neural networks. To improve the scalability of NAM, (chun-hao_chang_node-gam_2022) and (filip_radenovic_neural_2022) proposed the neural oblivious tree-based GAM and the basis network-based GAM, respectively. (xu_sparse_2023) introduced a sparse version of NAM using the group LASSO. One disadvantage of GAMs is their limited predictive power, which stems from the fact that they only learn first-order feature interactions – i.e., relationships between the target value and individual features. To address this, various studies have been conducted to enhance the predictive powers of GAMs by incorporating higher-order feature interactions, while still maintaining transparency. GA2M (lou_accurate_2013) simply takes pairwise features as inputs to learn pairwise interactions. GAMI-Net (zebin_yang_gami-net_2021), a neural network-based GAM, consists of networks for main effects (i.e., first-order interactions) and pairwise interactions. To enhance the interpretability of GAMI-Net, the sparsity and heredity constraints are added, and trivial features are pruned in the training process. Sparse interaction additive network (enouen_sparse_2022) is a 3-phase method for exploiting higher-order interactions. Initially, a black-box neural network is trained; subsequently, explainable feature attribution methods like LIME (ribeiro_why_2016) and SHAP (lundberg_unified_2017) identify the top-k important features, and finally, NAM is trained with these extracted features. (dubey_scalable_2022) introduced scalable polynomial additive model, an end-to-end model that learns higher-order interactions via polynomials. Similarly, (kim_higher-order_2022) proposed higher-order NAM that utilizes the feature crossing technique to capture higher-order interactions. Despite their capabilities, the aforementioned GAMs cannot process time series data, which limits their applicability in real-world scenarios. Recently, neural additive time series model (NATM) (jo_neural_2023), a time-series adaptation of NAM, has been proposed. However, NATM handles each time step independently with separate feature networks. This approach cannot capture temporal patterns and only takes a fixed-length time series as input. Our GATSM not only captures temporal patterns but also handles dynamic-length time series.
Appendix D Dataset details
@@
We use eight publicly available datasets for our experiments. Three datasets - Energy, Rainfall, and AirQuality - can be downloaded from the Monash repository (tan_monash_2020). Another three datasets - Heartbeat, LSST, and NATOPS - are available from the UCR repository (bagnall_uea_2018). The remaining two datasets can be downloaded from the PhysioNet (goldberger_physiobank_2000). Table 8 shows the statistics of the experimental datasets. Details of the datasets are provided below:
@@
  • Energy (chang_wei_tan_2020_3902637): This dataset consists of 24 features related to temperature and humidity from sensors and weather conditions. These features are measured every 10 minutes. The goal of this dataset is to predict total energy usage.
@@
  • Rainfall (chang_wei_tan_2020_3902654): This dataset consists of temperatures measured hourly. The goal of this dataset is to predict total daily rainfall in Australia.
@@
  • AirQuality (chang_wei_tan_2020_3902667): This dataset consists of features related to air pollutants and meteorological data. The goal of this dataset is to predict the PM10 level in Beijing.
@@
  • Heartbeat (liu_classification_2016): This dataset consists of heart sounds collected from various locations on the body. Each sound was truncated to five seconds, and a spectrogram of each instance was created with a window size of 0.061 seconds with a 70% overlap. The goal of this dataset is to classify the sounds as either normal or abnormal.
@@
  • Mortality (silva_predicting_2012) This dataset consists of records of adult patients admitted to the ICU. The input features include the patient demographics, vital signs, and lab results. The goal of this dataset is to predict the in-hospital death of patients.
@@
  • Sepsis (reyna_early_2019): This dataset consists of records of ICU patients. The input features include patient demographics, vital signs, and lab results. The goal of this dataset is to predict sepsis six hours in advance at every time step.
@@
  • LSST (emille_plasticc_2018): This challenge dataset aims to classify astronomical time series. These time series consist of six different light curves, simulated based on the data expected from the Large Synoptic Survey Telescope (LSST).
@@
  • NATOPS (alonso_aaltd_2016): This dataset aims to classify the Naval Air Training and Operating Procedures Standardization (NATOPS) motions used to control aircraft movements. It consists of 24 features representing the x, y, and z coordinates for each of the eight sensor locations attached to the body.
@@
We used get_UCR_data() and get_Monash_regression_data() functions in the tsai=0.3.9 library (tsai) to load the UCR and Monash datasets.
Appendix E Implementation details
@@
Table 9: Optimal hyper-parameters for GATSM.
We use 13 models, including GATSM, for our experiments. We implement XGBoost and EBM using the xgboost=2.0.3 (chen_xgboost_2016) and interpretml=0.6.1 (nori_interpretml_2019) libraries, respectively. For NodeGAM, we employ the official implementation provided by its authors (chun-hao_chang_node-gam_2022). We developed the remaining models using torch=2.0.1 (paszke_pytorch_2019). In addition, we implement the feature functions in NAM and NBM using grouped convolutions (krizhevsky_imagenet_2012; xie_aggregated_2017) to enhance their efficiency. We trained XGBoost and EBM on two AMD EPYC 7513 CPUs, while we trained the other models on an NVIDIA A100 GPU with 80GB VRAM. All models undergo hyperparameter tuning via optuna=3.6.1 (akiba_optuna_2019) with the Tree-structured Parzen Estimator (TPE) algorithm (bergstra_algorithms_2011) in 100 trials. The hyperparameter search space and the optimal hyperparameters for the models are provided below:
@@
  • GATSM: We tune nbm_batchnorm in the descret set {False, True}, nbm_dropout in the continuous interval [0, 0.9], attn_emb_size in the integer interval [8, 128], attn_n_heads in the integer interval [1, 8], attn_dropout in the continuous interval [0, 0.9], learning_rate in the continuous interval [1e-3, 1e-2], and weight_decay in the continuous interval [1e-6, 1e-1] on a log scale. Table 9 provides the optimal hyper-parameters for GATSM across all experimental datasets.
Appendix F Additional experiments
Multi-step forecasting
@@ Appendix F Additional experiments
Table 10: Statistics of the forecasting datasets.
@@
Table 11: Predictive performance comparison on the forecasting datasets.
Multi-step forecasting is a crucial task in time series. To validate the performance of GATSM on the multi-step forecasting task, we conducted experiments using two sophisticated forecasting models, DLinear(zeng_transformers_2023) and PatchTST(nie_time_2023), on three datasets: Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, Traffic [^2] http://pems.dot.ca.gov, and Weather [^3] https://www.bgc-jena.mpg.de/wetter/. Table 10 provides statistics on the experimental datasets. The models use 672 hours (four weeks) of data as input to generate 24, 48, and 72 hours-ahead predictions.
Table 11 presents the mean average percentage errors (MAPEs) and standard deviations of the experimental models over five different random seeds. Overall, PatchTST demonstrates the highest predictive accuracy, while GATSM shows lower predictive performance compared to the two state-of-the-art black-box models, DLinear and PatchTST. However, our GATSM is a transparent model and offers a unique advantage in terms of interpretability.
Model Cancer($R^{2}\uparrow$)
DLinear 0.956($\pm$0.009)
PatchTST 0.953($\pm$0.006)
GATSM 0.906($\pm$0.016)
@@
We conducted an experiment using the synthetic tumor growth dataset developed by (geng2017prediction), which simulates temporal changes in tumor size. This dataset is beneficial for verifying that the trained GATSM appropriately captures the influence of input features. Tumor size at any given time step is determined by three factors: prior tumor size, chemotherapy, and radiotherapy. For this experiment, we set the coefficient of chemotherapy to 10 and the coefficient of radiotherapy to 1; that is, chemotherapy more significantly reduces tumor size than radiotherapy. The experimental results in Table 12 showed that although GATSM underperforms compared to the complex black-box forecasting models, it nonetheless achieves strong predictive accuracy ($R^{2}>0.9$). Figure 4 and 4 illustrate the importance of time steps, and the contributions of the three factors. The importance highlights that recent time steps are notably more influential than earlier time steps due to the direct impact of prior tumor size. GATSM also effectively captured the effects of chemotherapy and radiotherapy on reducing tumor size, with chemotherapy having a significantly greater impact than radiotherapy (the scale of the y-axis in chemotherapy is greater than radiotherapy).
Appendix G Visualizations
In this section, we provide the visualizations of GATSM’s predictions for the AirQuality, Rainfall, and Mortality datasets.
AirQuality
Time-step importance
We plot the average attention scores at the last time step $T$ in Figure 5. The process for extracting the average attention score of time step $u$ at time step $t$ is formalized as $\sum_{k=1}^{K}a_{k,t,u}$. This process is repeated over all data samples, and the results are averaged. Based on Figure 5, it seems that GATSM pays more attention to the initial and last states than to the intermediate states. This indicates that the current concentration of particulate matter depends on the initial and recent states.
Global feature contribution
Figure 6 illustrates the global behavior of features in the AirQuality dataset, with red bars indicating the density of training samples. We extract $\sum_{k=1}^{K}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ from GATSM and repeat this process over the range of minimum to maximum feature values to plot the line. We found that the behavior of SO2, O3, and windspeed is inconsistent with prior human knowledge. Typically, high levels of SO2 and O3 are associated with poor air quality. However, GATSM learned that particulate matter concentration starts to decrease when SO2 exceeds 10 and O3 exceeds 5. This discrepancy may be due to sparse training samples in these regions, leading to insufficient training, or there may be interactions with other features. Another known fact is that high windspeed decreases particulate matter concentration. This is consistent when windspeed is below 0.7 in our observation. However, particulate matter concentration drastically increases when windspeed exceeds 0.7, likely due to the wind causing yellow dust.
Local time-independent feature contribution
To interpret the prediction of a data sample, we plot the local time-independent feature contributions, $\sum_{k=1}^{K}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$, in Figure 7. The main x-axis (blue) represents feature contribution, the sub x-axis (red) represents feature value, and the y-axis represents time steps. We found that SO2, NO2, CO, and O3 have positive correlations. In contrast, temperature, pressure, dew point, and windspeed have negative correlations. These are consistent with the global interpretations shown in Figure 6. Rainfall has the same values across all time steps.
Local time-dependent feature contribution
We also visualize the local time-dependent feature contributions, $\sum_{k=1}^{K}a_{k,t,u}h_{b}\left(x_{t,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$. Figure 8 illustrates the interpretation of the same data sample as in Figure 7. The time-dependent interpretation differs slightly from the time-independent interpretation. We found that there are time lags in SO2, NO2, CO, and O3, meaning previous feature values affect current feature contributions. For example, in the case of SO2, low feature values around time step 5 lead to low feature contributions around time step 13.
Rainfall
Time-step importance